Because of Erwin Chargaff's experiments on the proportion of DNA nucleotides, what percentage of a DNA molecule would be adenine if you had 30% cytosine?

Group of answer choices

A. 20%

B. 40%

C. 30%

D. 10%

2. Which of the following terms describes all of the genetic material of an organism?

Group of answer choices

A. Genome

B. Chromosome

C. Gene

D. Translation
Answer #1

Q1 Option A - 20%

G=C = 30%

There fore 100 - Percentage of (G+C) = A+T

G+C = 30+30 = 60%

100-60 = 40%

A+T = 40%

There fore A = 20%.

Q2.

Option A - Genome.

Genome represents all the genes, introns, non-coding regions, repetitive parts, entire DNA in the organism.